Two resistors, 35.5 and 82.3Ω, are connected in parallel. The current through the 82.3-Ω resistor is 4.13A. (a) Determine the current in the other resistor. (b) What is the total power supplied to the two resistors?
1
Expert's answer
2020-09-08T09:11:04-0400

a) If the two resistors are connected in parallel, the voltage in the resistors will be equal. We may calculate the voltage in the second resistor:

"V=R_2 I_2= 82.3\cdot4.13 \approx 339.9" volts.

The same voltage is in the first resistor, so the current in it can be obtained as

"I_1 = \dfrac{V}{R_1} = \dfrac{339.9}{35.5} \approx 9.57" A.

b) The total power can be obtained by taking into account the voltage and the total current "I=I_1+I_2."

The total power is "P=I\cdot V = (I_1+I_2)V = (9.57+4.13)\cdot339.9\approx 4657 W."